OpenStudy (anonymous):

Indicate the equation of the line that is the perpendicular bisector of the segment with endpoints (4, 1) and (2, -5).

OpenStudy (anonymous):



OpenStudy (campbell_st):

start by finding the slope of the line joining the 2 points... and also find the midpoint of the segment

OpenStudy (anonymous):

The slope is 3 right?

OpenStudy (anonymous):

And the midpoints are 3, -2 right?

OpenStudy (campbell_st):

slope is correct... and midpoint is correct. so you need to find the negative reciprocal of the slope because the product the the slope of a line and a perpendicular must be -1 so the slope of the perpendicular is m = -1/3 so the equation of the line is \[y = - \frac{1}{3} x+ b\] just substitute the x and y values from the midpoint to find b. hope it helps

OpenStudy (anonymous):

I dont get it :( I need to get the answer without any fractions

OpenStudy (campbell_st):

well then multiply every term in the equation by the denominator of the fraction... so what value did you get for b..?

OpenStudy (anonymous):

how do I get b??

OpenStudy (campbell_st):

your midpoint is (3, -2) so x = 3 and y = -2 substitute them into \[y = -\frac{1}{3} x + b\] what do you get..?

OpenStudy (anonymous):

is it -1?

OpenStudy (campbell_st):

well that's correct so the equation is \[y = -\frac{1}{3} x - 1\] if you don't like the fraction multiply each term by 3

OpenStudy (campbell_st):

the equation can be 3y = -x - 3 or x + 3y = -3

OpenStudy (anonymous):

if I multiply by 3..what would I get? would it be x+y=-3?

OpenStudy (anonymous):

but it wont fit my answer table.